### v1

@@ 3.3 Measured results
Figure 5a illustrates the test setup employed to validate the BFIC and the proposed dynamic deformation correction technique. The wideband modulated signals are generated in MATLAB and transmitted through a Pasternack PE9887 horn antenna using DAC on Xilinx ZCU216 RFSoC. The 3D-printed array, along with the BFIC, is mounted on the DAMS D6025 antenna measurement platform as shown in Fig. 5b. A 4.2GHz differential clock is provided from ADF4372 EVM (wideband synthesizer) to the internal LO and sampler clock generation circuit of the BFIC. The bias voltages for both the BFIC and the self-calibration loop are provided using a DAC81416 EVM (DAC) and the serial control bits to the on-chip serial to parallel interface (SPI) registers being applied from a Digilent Analog Discovery 2 board. The quadrature $BF_{\text{out}}$ outputs post combining from BFIC is captured on the Xilinx ZCU216 ADCs and subsequently converted to a 16-bit digital word using a mapping function. The digital $BF_{\text{out}}$ is then fed to the calibration loop, interfaced through a Waveshare QFN-64 programmable adapter socket. After each iteration, the new control bits for each phase are captured from the respective calibration loops and updated to the serial control bit stream. This is then fed to the SPI control registers using Digilent Analog Discovery II. The entire data capture loop is automated through the power automation tool .
To assess the worst-case bending and verify the effectiveness of the calibration technique, a $4\times 4$ array is deformed to a curvature with a radius of 38cm using a custom-designed antenna holder as shown in Fig. 5e. Within this array, three tiles are disabled, leaving only a $2\times 2$ tile active under deformation, resulting in a maximum beam pointing error of $7^{\circ}$. One antenna is kept as reference (REF) while the calibration loop is connected to the other three antennas to optimize the corresponding PS $\phi_{1}$, $\phi_{2}$, $\phi_{3}$. Fig. 5c shows phase control codes are auto-adjusting to achieve the optimal beamformed output under this deformation. Each loop is settled to minimize the beam pointing error $<1.5^{\circ}$. This error can be further reduced by increasing the PS resolution. The 2dB gain reduction after beam-point error correction is due to the antenna holder and the change in Line-of-Sight (LOS). Figure 5d shows the radiation pattern without deformation and with deformation after self-correction. The FoV enhancement after deformation is evident from the radiation pattern.
@@
Figure 6: a. The chip micrograph shows the proposed conformal array beamforming receiver, along with the calibration loop. The circular patch antenna,(bottom right) is printed on Ninjaflex using silver ink. The receiver IC and its connections, (top left) are printed using Cu ink on an AP layer.b.Additively Printed $2\times 2$ conformal array top and back sides. c. Azimuth and d. Elevation patterns of the $2\times 2$ conformal array. e.S11 and conversion gain of the single receiver channel
@@ 6 AUTHOR CONTRIBUTIONS
This project was conceived by S.G., S.P., J.N., and J.W. The antenna was designed and additively printed by T.D. and K.K with inputs from P.L.. The Cu ink was developed and verified by A. I. and S.R. The receiver front-end was designed by S.P., A.V., Q.X., E.G., and A.R. with inputs from S.G. and S.S. The discrete-time combiner was designed by E.G., S.P., and A.R. with inputs from S.G.. The phase-interpolator and time-interleaved clock generator was designed by S.P., and Q.X. with inputs from S.S. and S.G.. Calibration loop, design, and validation was done by S.B. and S.P. with inputs from S.G. Measurements of the self-calibrated array is performed by S.P. with inputs from S.G.. S.P., A.R., and S.G. are equal contributors for preparing the manuscript.
@@ 8.2 Single tile $2\times 2$ Receiver Array Performance
Figure 7: Measured beamforming RX performance including a. Beamforming gain over wideband modulated input with 1-element and 4-element b. EVM for 16-QAM modulation c.Out-Of-Band IIP3 d.In-Band IIP3 e.i-dB compression point.
@@
Figure 8: a.Weight measurement of $2\times 2$ tile with BFIC b. Power breakdown of the conformal tile
Figure 8b shows the power breakdown for the proposed BFIC with the four RF front-end units consuming highest power. The LNTA, passive mixer, TIA, and buffer together with few test circuits consume  41mW in the total power. Clock drivers, LO Generation, LO PS, sampler clock generation together with the PI and the time-interleaved sample-and-hold clocking consumes 27 mW. The charge-domain summer which includes a 2-stage internally compensated OTA consumes 6 mW including both the quadrature paths. The SPI registers only consume 26 $\mu$W. The 3 variable calibration loop consumes 4.5 mW.

### v3

@@ 3.3 Measured results
Figure 5a illustrates the test setup employed to validate the BFIC and the proposed dynamic deformation correction technique. The wideband modulated signals are generated in MATLAB and transmitted through a Pasternack PE9887 horn antenna using DAC on Xilinx ZCU216 RFSoC. The 3D-printed array, along with the BFIC, is mounted on the DAMS D6025 antenna measurement platform as shown in Fig. 5b. A 4.2GHz differential clock is provided from ADF4372 EVM (wideband synthesizer) to the internal LO and sampler clock generation circuit of the BFIC. The bias voltages for both the BFIC and the self-calibration loop are provided using a DAC81416 EVM (DAC) and the serial control bits to the on-chip serial to parallel interface (SPI) registers being applied from a Digilent Analog Discovery 2 board. The quadrature $BF_{\text{out}}$ outputs post combining from BFIC is captured on the Xilinx ZCU216 ADCs and subsequently converted to a 16-bit digital word using a mapping function. The digital $BF_{\text{out}}$ is then fed to the calibration loop, interfaced through a Waveshare QFN-64 programmable adapter socket. After each iteration, the new control bits for each phase are captured from the respective calibration loops and updated to the serial control bit stream. This is then fed to the SPI control registers using Digilent Analog Discovery II. The entire data capture loop is automated through the power automation tool [50].
To assess the worst-case bending and verify the effectiveness of the calibration technique, a $4\times 4$ array is deformed to a curvature with a radius of 38cm using a custom-designed antenna holder as shown in Fig. 5e. Within this array, three tiles are disabled, leaving only a $2\times 2$ tile active under deformation, resulting in a maximum beam pointing error of $7^{\circ}$. One antenna is kept as reference (REF) while the calibration loop is connected to the other three antennas to optimize the corresponding PS $\phi_{1}$, $\phi_{2}$, $\phi_{3}$. Fig. 5c shows phase control codes are auto-adjusting to achieve the optimal beamformed output under this deformation. Each loop is settled to minimize the beam pointing error $<1.5^{\circ}$. This error can be further reduced by increasing the PS resolution. The 2dB gain reduction after beam-point error correction is due to the antenna holder and the change in Line-of-Sight (LOS). The phase control codes are properly initialized so that the calibration loop doesn’t get stuck at side-lobes (local maxima). Figure 5d shows the radiation pattern without deformation and with deformation after self-correction. The FoV enhancement after deformation is evident from the radiation pattern.
@@
Figure 6: a. The $2\times 2$ conformal array tile back side, the receiver IC and its connections, are printed using Cu ink on an AP layer, chip micrograph of beamforming receiver, along with the calibration loop. b. Circular patch antenna printed on Ninjaflex using silver ink. c. Azimuth and d. Elevation patterns of the $2\times 2$ conformal phased array. e. S11 and conversion gain of the single receiver channel.
@@ 6 AUTHOR CONTRIBUTIONS
This project was conceived by S.G., S.P., J.N., and J.W. The antenna was designed and additively printed by T.D. and K.K with inputs from P.L.. The Cu ink was developed and verified by A. I. and S.R. The receiver front-end was designed by S.P., A.M., Q.X., E.G., and A.R. with inputs from S.G. and S.S. The discrete-time combiner was designed by E.G., S.P., and A.R. with inputs from S.G.. The phase-interpolator and time-interleaved clock generator was designed by S.P., and Q.X. with inputs from S.S. and S.G.. Calibration loop, design, and validation was done by S.B. and S.P. with inputs from S.G. Measurements of the self-calibrated array were performed by S.P. with inputs from S.G.. S.P., A.R., and S.G. are equal contributors for preparing the manuscript.
@@ 8.2 Single tile $2\times 2$ Receiver Array Performance
Figure 7: Measured beamforming RX performance including a. Beamforming gain over wideband modulated input with 1-element and 4-element b. EVM for 16-QAM modulation c. out-Of-band IIP3 d. in-band IIP3 e.1-dB compression point.
@@
The power breakdown for the proposed BFIC is given below, the four RF front-end units consuming highest power. The LNTA, passive mixer, TIA, and buffer together with few test circuits consume  41mW in the total power. Clock drivers, LO Generation, LO PS, sampler clock generation together with the PI and the time-interleaved sample-and-hold clocking consumes 27 mW. The charge-domain summer which includes a 2-stage internally compensated OTA consumes 6 mW including both the quadrature paths. The SPI registers only consume 26 $\mu$W. The 3 variable calibration loop consumes 4.5 mW.